OpenStudy (anonymous):

What's the product (-2 1/2) (-3 1/5)

hero (hero):

Hints: \(- \times - = + \) 1/2 = .5 1/5 = .2 Therefore, (-2 1/2) (-3 1/5) = 2.5 x 3.2

hero (hero):

Or you can mutliply \((2 + \frac{1}{2})(3 + \frac{1}{5})\)

OpenStudy (anonymous):

So the answers 8